OpenStudy (anonymous):

Kara's favorite jeans are on sale 30% off. They retail for $70.00 How much will she save? A. $14.00 B. $17.00 C. $21.00 D. $30.00

OpenStudy (anonymous):

If you take .30 times 70.00 what will you get? This will help you with your answer

OpenStudy (anonymous):

70*0.30=$21 saved. 70-21=$49 is what she'll spend.

OpenStudy (anonymous):

.3*70=21 C is right

OpenStudy (anonymous):

Yes, taht is correct